OpenStudy (anonymous):

make a the subject of the formula s=a/4+8u will medal

OpenStudy (anonymous):

@UsukiDoll

undeadknight26 (undeadknight26):

s = a/4 + 8u Subtract both sides by 8u. s - 8u = a/4 Multiply both sides by 4. a = 4(s - 8u)

OpenStudy (usukidoll):

are we solving for a in this question?

undeadknight26 (undeadknight26):

O.O

OpenStudy (usukidoll):

I just want clarification before I go further

OpenStudy (usukidoll):

@Sophhh786 is this question about solving for a ? as in having a by itself?

OpenStudy (anonymous):

yes

OpenStudy (usukidoll):

alright so given the equation \[s= \frac{a}{4}+8u\] is this your equation?

OpenStudy (anonymous):

yes

OpenStudy (usukidoll):

alright so.. we need to get rid of the fraction. Therefore multiply 4 throughout the equation. Can you do that?

OpenStudy (anonymous):

yes 4s =a+8u

OpenStudy (usukidoll):

close. the 4s and the a is right but for the 8u it should be 4(8u) = 32u

OpenStudy (anonymous):

why?

OpenStudy (usukidoll):

because we are multiplying 4 throughout the equation. every term is going to be multiplied by 4

OpenStudy (usukidoll):

\[(4)s= \frac{a(4)}{4}+8u (4)\]

OpenStudy (anonymous):

so it is like linear equations

OpenStudy (usukidoll):

yeah

OpenStudy (usukidoll):

so now we have \[4s=a+32u\] so we just need one more step. What do we need to do to get a by itself?

OpenStudy (anonymous):

divide 4 y both sides

OpenStudy (anonymous):

by

OpenStudy (usukidoll):

not exactly.. we need a by itself and we have this equation. it's like get all the terms that's NOT a to the left so I need to subtract both sides by 32u

OpenStudy (anonymous):

i still dont understand why we need to multiply everything by 4 in the equation u should just multiply the 4 in the fraction and s ?

OpenStudy (usukidoll):

we multiplied 4 throughout the equation at the beginning to get rid of the fraction

OpenStudy (usukidoll):

there's more than one way to solve this... first is what I and undeadknight did earlier which was to multiply 4 throughout the equation to get rid of the fraction that was attached to a. then subtract 32u on both sides. it will look something like this \[4s=a+32u \] \[4s-32u=a+32u-32u \] \[4s-32u=a \]

OpenStudy (usukidoll):

so let's do an alternative version where we don't multiply 4 throughout the equation first.

OpenStudy (anonymous):

ok

OpenStudy (usukidoll):

\[s= \frac{a}{4}+8u \] is the original equation. and we need a by itself. so I have to shift all terms (except a because we need it by itself to the right) therefore subtract 8u on both sides

OpenStudy (usukidoll):

\[s-8u= \frac{a}{4}\] so we have a by itself... but now we have to divide each term by 1/4

OpenStudy (anonymous):

im gonna stick with the first method you told me im gonna go through it step by step now if I have any difficulties in understanding i will give you a shout yeah? Thanks though :)

OpenStudy (usukidoll):

ok. :) yeah I prefer what @undeadknight26 did... for dividing 1/4 all over the equation.. we have to flip that second fraction (it will become 4/1 and then you have to use multiplication) because there's no such thing as dividing fractions.

OpenStudy (anonymous):

yeh actually @undeadknight26 version makes sense to me

OpenStudy (usukidoll):

and fewer steps too :)

OpenStudy (anonymous):

the only thing i dont get is why has he inserted brackets has he factorized in the last stage or something

OpenStudy (usukidoll):

oh ... in that step there was a 4 in common so he took it out

OpenStudy (usukidoll):

a=4(s-8u) distributing the 4 gives us a=4s-32u now we split up the number 32 a=4s-(4)(8)u there is a 4 on each term so we yank it out a=4(s-8u)

OpenStudy (anonymous):

could you write it like this though a=4s-32 would that be appropriate for the final answer?

OpenStudy (usukidoll):

yeah

OpenStudy (anonymous):

ok thanks @undeadknight26 and @UsukiDoll

OpenStudy (usukidoll):

you're welcome :)